OpenStudy (anonymous):

Can someone please help me find this derivative I know the answer but I don't get how to find it : x (x^2 + 1) ^(-1/2)

OpenStudy (anonymous):

The step I am up to is 1 (x^2 + 1) ^ (-1/2) + x (-1/2) (x^2 + 1) ^-3/2 (2x) but I don't know where to go after this

OpenStudy (anonymous):

(x^2 + 1) ^ (-1/2) is common in both

OpenStudy (anonymous):

take it out and then add

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you elaborate further? add ?

OpenStudy (anonymous):

like can you show me

OpenStudy (anonymous):

(x^2 + 1) ^ (-1/2) [ 1 + x (-1/2) (x^2 + 1) ^-1 (2x) ]

OpenStudy (anonymous):

is this clear? :)

OpenStudy (anonymous):

ok the final answer was 1 / (x^2 + 1) ^ -3/2 how would u get that from this?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

? can i get an answer

OpenStudy (anonymous):

well...i am not getting that answer :|

OpenStudy (anonymous):

yeah that is the correct answer though

OpenStudy (anonymous):

can someone thats looking at this prob guide me through with steps?

OpenStudy (asnaseer):

Is this the equation you are trying to differentiate?\[x\sqrt{x^2+1}\]

OpenStudy (asnaseer):

sorry - I meant:\[\frac{x}{\sqrt{x^2+1}}\]

OpenStudy (anonymous):

yeah asnnaseer!

OpenStudy (anonymous):

i got it :D

OpenStudy (anonymous):

Can you guide me through from the beginning?

OpenStudy (anonymous):

please

OpenStudy (asnaseer):

ok - it looks like @telltoamit has got it so I'll let him/her explain it to you...

OpenStudy (anonymous):

(x^2 + 1) ^ (-1/2) [ 1 + x (-1/2) (x^2 + 1) ^-1 (2x) ] = (x^2 + 1) ^ (-1/2) [ 1 - x^2 (x^2 + 1) ^-1 ]

OpenStudy (anonymous):

(x^2 + 1) ^ (-1/2) [ (x^2 + 1) ^-1 ]

OpenStudy (anonymous):

thats ur answer

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

yeah thankss

OpenStudy (anonymous):

You're welcome :)